# VoxAct-B: Voxel-Based Acting and Stabilizing Policy for Bimanual Manipulation

### v1

@@ Abstract
Bimanual manipulation is critical to many robotics applications. In contrast to single-arm manipulation, bimanual manipulation tasks are challenging due to higher-dimensional action spaces. Prior works leverage large amounts of data and primitive actions to address this problem, but may suffer from sample inefficiency and limited generalization across various tasks. To this end, we propose VoxAct-B, a language-conditioned, voxel-based method that leverages Vision Language Models (VLMs) to prioritize key regions within the scene and reconstruct a voxel grid. We provide this voxel grid to our bimanual manipulation policy to learn acting and stabilizing actions. This approach enables more efficient policy learning from voxels and is generalizable to different tasks. In simulation, we show that VoxAct-B outperforms strong baselines on fine-grained bimanual manipulation tasks. Furthermore, we demonstrate VoxAct-B on real-world Open Drawer and Open Jar tasks using two UR5s. Code, data, and videos will be available at https://voxact-b.github.io.
@@ 1 Introduction
Bimanual manipulation is essential for robotics tasks, such as when objects are too large to be controlled by one gripper or when one arm stabilizes an object of interest to make it simpler for the other arm to manipulate [1]. In this work, we focus on asymmetric bimanual manipulation. Here, “asymmetry” refers to the functions of the two arms, where one is a stabilizing arm, while the other is the acting arm. Asymmetric tasks are common in household and industrial settings, such as cutting food, opening bottles, and packaging boxes. They typically require two-hand coordination and high-precision, fine-grained manipulation, which are challenging for current robotic manipulation systems. To tackle bimanual manipulation, some methods [2, 3] train policies on large datasets, and some exploit primitive actions [4, 5, 6, 7, 8, 9, 10]. However, they are generally sample inefficient, and using primitives can hinder generalization to different tasks.
To this end, we propose VoxAct-B, a novel voxel-based, language-conditioned method for bimanual manipulation. Voxel representations, when coupled with discretized action spaces, can increase sample efficiency and generalization by introducing spatial equivariance into a learned system, where transformations of the input lead to corresponding transformations of the output [11]. However, processing voxels is computationally demanding [12, 13]. To address this, we propose utilizing VLMs to focus on the most pertinent regions within the scene by cropping out less relevant regions. This substantially reduces the overall physical dimensions of the areas used to construct a voxel grid, enabling an increase in voxel resolution without incurring computational costs. To our knowledge, this is the first study to apply voxel representations in bimanual manipulation.
@@
We extend the RLBench [14] benchmark to support bimanual manipulation. We introduce an asymmetric bimanual version of Open Drawer, Open Jar, and Put Item in Drawer tasks. VoxAct-B outperforms strong baselines, such as ACT [3], Diffusion Policy [15], VoxPoser [16], and PerAct [11], by a large margin. We also validate our approach on a real-world bimanual manipulation setup with two UR5s on Open Drawer and Open Jar. See Figure 1 for an overview.
@@
Figure 1: VoxAct-B uses voxel representations and language to perform bimanual manipulation with 6-DoF manipulation from both arms. We test three language-conditioned bimanual tasks in simulation and two (Open Drawer and Open Jar) on a real-world setup with two UR5s. The prompt for Open Drawer assumes the left arm is stabilizing and the right arm is acting, while the reverse is true for the Open Jar prompt.
@@ 2 Related Work
Action Space Representation. For 2D manipulation, prior works have shown the benefits of action representations based on spatial action maps [21, 41, 42, 43, 44, 45, 46], including in bimanual contexts [10, 21], where neural networks directly predict 2D “images” that indicate desired locations for the action. Compared to directly regressing the action location, using spatial action maps better handles multimodality and has 2D equivariance, where translations and rotations of the input image map to similar transformations of the output action. Recent works have extended this idea to support 3D spatial action maps, which classify an action’s location as a 3D point in the robot’s workspace, and thus maintain spatial equivariance. For example, PerAct [11] is a language-conditioned behavioral cloning agent that takes voxel grids as input and outputs 6-DoF actions. While PerAct achieved state-of-the-art performance on RLBench, it has a high computational cost due to processing voxels. Follow-up works, such as RVT [12] and Act3D [13], have reduced the computational cost of PerAct by avoiding voxel representations but often need multiple views of the scene to achieve optimal performance and may be less interpretable compared to a voxel grid that contains a 3D spatial action map. These prior works have also not been applied to bimanual manipulation. In this work, we retain the spatial equivariance benefits of voxel representations but reduce the cost of processing voxels by “zooming” into part of the voxel grid. This is similar to the intent of C2F-ARM [47], but we use the knowledge in VLMs to determine the most relevant regions in the voxel grid.
@@ 3 Problem Statement
Given access to a pre-trained VLM and expert demonstrations, the objective is to produce a bimanual policy $\pi$ for a variety of language-conditioned manipulation tasks. We assume a flat workspace with two fixed-base robot manipulators, each with a parallel-jaw gripper. A policy $\pi$ controls both arms by producing actions $a_{t}=(a_{t}^{s},a_{t}^{a})$ at each time step $t$, where $a_{t}^{s}$ and $a_{t}^{a}$ follow [39] and refer to the stabilizing and acting arm actions, respectively. For simplicity, we suppress the time $t$ when the distinction is unnecessary. We use the low-level action representation from PerAct [11] with $a^{s}=(a_{\rm pose}^{s},a_{\rm open}^{s},a_{\rm collide}^{s})$ and $a^{a}=(a_{\rm pose}^{a},a_{\rm open}^{a},a_{\rm collide}^{a})$. These specify each arm’s 6-DOF gripper pose, its gripper open state, and whether a motion planner for the arms used collision avoidance to reach an intermediate pose. We assume task-specific demonstrations $\mathcal{D}_{\mathcal{\ell}}=\{\zeta_{1},\zeta_{2},\ldots,\zeta_{n}\}$ and two common language commands $\ell_{as}$ and $\ell_{sa}$, where $as$ denotes the left arm as acting and right arm as stabilizing, and vice versa for $sa$. Each demonstration consists of a set of keyframes extracted from a sequence of continuous actions paired with observations. We adapt the keyframe extraction function from [11] by including keyframes that have an action with near-zero joint velocities and unchanged gripper open state for acting and stabilizing arms. The observation at each time is the 3D voxel grid $\mathbf{v}$ of dimension $(L\times W\times H)$, where we use $\mathbf{v}[x,y,z]$ to denote an individual voxel at coordinates $(x,y,z)$. The robot also receives the language command $\mathbf{l}\in\{\ell_{as},\ell_{sa}\}$, which is fixed for all time steps in an episode, where the robot interacts with the environment for up to $T$ time steps. An episode terminates with a task-dependent success criteria or failure (if otherwise).
@@ 4.1 Extending PerAct for Bimanual Manipulation
PerAct [11] was originally designed and tested for single-arm manipulation. We extend it to support bimanual manipulation. A natural way to do this would be to train separate policies for the two arms. However, we exploit the discretized action space that predicts the next best voxel with spatial equivariance properties and formulate a system that uses acting and stabilizing policies. In contrast to a policy that operates in joint-space control, acting and stabilizing policies perform the same functions irrespective of whether it is a left arm or a right arm, assuming the next best voxel is kinematically feasible for both arms. Hence, either arm can execute an acting policy or a stabilizing policy, which improves policy learning. In the low-level action space, the arms execute one low-level action $a_{t}^{s}$ and $a_{t}^{a}$ (see Section 3) at each time $t$. In the following, we use similar notation as [11] but index components as belonging to an arm using the superscript: $\textrm{arm}\in\{\textrm{acting},\textrm{stabilizing}\}$.
At each time step, the input to each arm is a voxel observation $\mathbf{v}$, proprioception data of both robot arms $\rho$, a language goal $\mathbf{l}\in\{\ell_{as},\ell_{sa}\}$, and an arm ID $\xi\in\{0,1\}$, and the task is to predict an action. During training, the language goal is given in the data, but during evaluation, we use VLMs to determine which language goal, $\ell_{as}$ or $\ell_{sa}$, to use. If the language goal is $\ell_{as}$, we assign the left arm ($\xi=0$) to the acting policy and the right arm ($\xi=1$) to the stabilizing policy, and conversely for $\ell_{sa}$. During training, this allows our method to learn to map the appropriate acting or stabilizing actions to a given arm, and during evaluation, this informs each arm’s actions. Note that the predicted arm ID is discarded. PerAct uses value maps to represent different components of the action space, where predictions for each arm are $\mathcal{Q}$-functions with state-action values. Formally, we have the following five value maps per arm, as the output of the arm’s learned deep neural network, where:
@@
The demonstrations provide labels for each arm’s five action components, giving us the following nine label sources: ${Y_{\textrm{trans}}^{\textrm{arm}}\in\mathbb{R}^{L\times W\times H}}$ for translations, ${Y_{\textrm{rot}}^{\textrm{arm}}\in\mathbb{R}^{(360/R)\times 3}}$ (with $R=5$) for discretized rotations, ${Y_{\textrm{open}}^{\textrm{arm}}\in\mathbb{R}^{2}}$ for the binary open variables, ${Y_{\textrm{collide}}^{\textrm{arm}}\in\mathbb{R}^{2}}$ for the binary collide variables, and ${Y_{\textrm{id}}^{\textrm{arm}}\in\mathbb{R}^{2}}$ for the binary arm ID variables. The overall training loss for VoxAct-B is:
@@ 4.2 VoxAct-B: Voxel Representations and PerAct for Bimanual Manipulation
To detect the object of interest reliably, we use a two-stage approach similar to [16]. We input a text query to an open-vocabulary object detector to detect the object. Then, we use a foundational image segmentation model to obtain the segmentation mask of the object and use the mask’s centroid along with point cloud data to retrieve the object’s pose with respect to the front camera. We use the pose of the object to determine the task-specific roles of each arm and the language goal. This cropped voxel grid and language goal are the input to our bimanual manipulation policy. We call our method VoxAct-B: Voxel-Based Acting and Stabilizing Policy. See Figures 2 and 3 for an overview.
4.3 Additional Implementation Details
[…]
@@ 5 Experiments
5.1 Tasks
In simulation, we build on top of RLBench [14], a popular robot manipulation benchmark widely used in prior work, including VoxPoser and PerAct. We extend it to support bimanual manipulation (see Appendix A for details). We do not perform simulation-to-real transfer in this paper; simulation is for algorithm development and benchmarking. We design the following three bimanual tasks:
  • Open Jar: a jar with a screw-on lid is randomly spawned and scaled from 90% to 100% of the original size within the robot’s $0.43\times 0.48$ meters of workspace. The jar color is uniformly sampled from a set of 20 colors. The robot must first grasp the jar with one hand and use the other to unscrew the lid in an anti-clockwise direction until it is completely removed.
@@ 5.1 Tasks
See Figure 1 for an illustration. In the real world, we test Open Jar and Open Drawer using a coffee jar with dimensions $3.35\times 2.85\times 4.8$ inches and a drawer of dimensions $12\times 12\times 12$ inches.
@@ 5.2 Baselines and Ablations
  • VoxAct-B w/o arm ID: excludes arm ID as input and disables the corresponding loss function.
@@ 5.3 Experiment Protocol and Evaluation
To generate demonstrations in simulation, we follow the convention from RLBench and define a sequence of waypoints to complete the task, and use motion planning to control the robot arms to reach waypoints. We generate 10 and 100 demonstrations of training data. Half of this data consists of left-acting and right-stabilizing demonstrations, and the other half contains right-acting and left-stabilizing demonstrations. We generate 25 episodes of validation and test data using different random seeds. We train and evaluate all methods using three random seeds and report the average of the results. We evaluate all methods on the same set of test demonstrations for a fair comparison.
Each method saves a checkpoint every 10,000 training steps. For all methods, we use the best-performing checkpoint, evaluated on the validation data, to obtain the test success rate and report this result. Deciding the best checkpoints for VoxAct-B and ablations is nontrivial since iterating over all possible combinations is computationally expensive. For example, with 400,000 training steps, using the same 10,000 checkpoint interval means there are $40\times 40=1600$ possible combinations. Therefore, with the validation data, we use the latest stabilizing checkpoint to evaluate all acting checkpoints; we use the best acting checkpoint to evaluate all stabilizing checkpoints. Then, we use the best-performing acting and stabilizing checkpoints to obtain the test success rate.
@@ 6.1 Simulation Results
Open Open Put Item
Jar Drawer in Drawer
Method 10 100 10 100 10 100
Diffusion Policy 5.3 24.0 4.0 6.7 2.7 6.7
ACT w/Transformers 1.3 26.7 10.7 29.3 8.0 50.7
VoxPoser 8.0 8.0 32.0 32.0 4.0 4.0
Bimanual PerActs 9.3 - 40.0 - 6.7 -
VoxAct-B (ours) 38.6 58.7 73.3 73.3 36.0 46.7
Table 1: Performance of different methods on bimanual manipulation tasks in simulation, based on 10 or 100 (task-specific) training demonstrations. We use three training seeds for all methods, and evaluate on the same 25 episodes of unseen test data using the best checkpoints from validation (Section 5.3). The results are the average evaluation over three seeds. We only test Bimanual Peracts with ten demonstrations (not 100) due to computational constraints. VoxPoser does not have training, so its 10 and 100 results are identical.
Comparisons with baselines. Table 1 reports the test success rates of baselines and VoxAct-B. When we train all methods using ten demonstrations, VoxAct-B outperforms all baselines by a large margin. In a low-data regime, the discretized action space with spatial equivariance properties (as used in VoxAct-B and Bimanual PerActs) may be more sample-efficient and easier for learning-based methods compared to methods that use joint space (ACT and Diffusion Policy). When we train all methods using more demonstrations (100), VoxAct-B still outperforms baselines on most tasks, except ACT for Put Item in Drawer. Through ablations of ACT and Diffusion Policy, we found that fixing the roles of acting and stabilizing arms greatly improved their performance. We theorize that they lack a grounding mechanism that allows both arms to learn acting and stabilizing actions effectively. We attribute the tasks’ difficulty to the following: high environment variation, difficult bimanual manipulation tasks with high-dimensional action spaces and fine-grained manipulation, and the two types of training data that each method needs to learn (based on which arms are acting and stabilizing).
@@
VoxAct-B w/o VLMs 16.0
VoxAct-B w/o acting and stabilizing 66.7
@@
Ablation experiments. Table 2 reports results on Open Drawer in simulation, based on 10 training demonstrations. We use the same training and evaluation protocols as Table 1. VoxAct-B w/o VLMs performs poorly versus VoxAct-B because, without using the VLMs to reduce the physical space, the voxel resolution is lower due to the large workspace area for each individual voxel, which hinders fine-grained manipulation. Moreover, VoxAct-B w/o acting and stabilizing and VoxAct-B w/o arm ID perform worse than VoxAct-B, and they also struggle with the same issues as the baselines.

### v2

@@ Abstract
Bimanual manipulation is critical to many robotics applications. In contrast to single-arm manipulation, bimanual manipulation tasks are challenging due to higher-dimensional action spaces. Prior works leverage large amounts of data and primitive actions to address this problem, but may suffer from sample inefficiency and limited generalization across various tasks. To this end, we propose VoxAct-B, a language-conditioned, voxel-based method that leverages Vision Language Models (VLMs) to prioritize key regions within the scene and reconstruct a voxel grid. We provide this voxel grid to our bimanual manipulation policy to learn acting and stabilizing actions. This approach enables more efficient policy learning from voxels and is generalizable to different tasks. In simulation, we show that VoxAct-B outperforms strong baselines on fine-grained bimanual manipulation tasks. Furthermore, we demonstrate VoxAct-B on real-world Open Drawer and Open Jar tasks using two UR5s. Code, data, and videos are available at https://voxact-b.github.io.
@@ 1 Introduction
Bimanual manipulation is essential for robotics tasks, such as when objects are too large to be controlled by one gripper or when one arm stabilizes an object of interest to make it simpler for the other arm to manipulate [1]. In this work, we focus on asymmetric bimanual manipulation. Here, “asymmetry” refers to the functions of the two arms, where one is a stabilizing arm, while the other is the acting arm. Asymmetric tasks are common in household and industrial settings, such as cutting food, opening bottles, and packaging boxes. They typically require two-hand coordination and high-precision, fine-grained manipulation, which are challenging for current robotic manipulation systems. To tackle bimanual manipulation, some methods [2, 3] train policies on large datasets, and some exploit primitive actions [4, 5, 6, 7, 8, 9, 10]. However, they are generally sample inefficient, and using primitives can hinder generalization to different tasks as they are not easily adaptable to other types of tasks.
To this end, we propose VoxAct-B, a novel voxel-based, language-conditioned method for bimanual manipulation. Voxel representations, when coupled with discretized action spaces, can increase sample efficiency and generalization by introducing spatial equivariance into a learned system, where transformations of the input lead to corresponding transformations of the output [11]. However, processing voxels is computationally demanding [12, 13]. To address this, we propose utilizing VLMs to focus on the most pertinent regions within the scene by cropping out less relevant regions. This substantially reduces the overall physical dimensions of the areas used to construct a voxel grid, enabling an increase in voxel resolution without incurring computational costs. To our knowledge, this is the first study to apply voxel representations in bimanual manipulation. Additionally, VoxAct-B does not rely on action primitives, making it more general and applicable to a wider range of tasks.
@@
We extend the RLBench [14] benchmark to support bimanual manipulation. We introduce a bimanual version of Open Drawer, Open Jar, Put Item in Drawer, and Hand Over Item tasks. VoxAct-B outperforms strong baselines, such as ACT [3], Diffusion Policy [15], VoxPoser [16], and PerAct [11], by a large margin. We also validate our approach on a real-world bimanual manipulation setup with two UR5s on Open Drawer and Open Jar. See Figure 1 for an overview.
@@
Figure 1: VoxAct-B uses voxel representations and language to perform bimanual manipulation with 6-DoF manipulation from both arms. We test four language-conditioned bimanual tasks in simulation and two (Open Drawer and Open Jar) on a real-world setup with two UR5s.
@@ 2 Related Work
Action Space Representation. For 2D manipulation, prior works have shown the benefits of action representations based on spatial action maps [21, 41, 42, 43, 44, 45, 46], including in bimanual contexts [10, 21], where neural networks directly predict 2D “images” that indicate desired locations for the action. Compared to directly regressing the action location, using spatial action maps better handles multimodality and has 2D equivariance, where translations and rotations of the input image map to similar transformations of the output action. Recent works have extended this idea to support 3D spatial action maps, which classify an action’s location as a 3D point in the robot’s workspace, and thus maintain spatial equivariance. For example, PerAct [11] is a language-conditioned behavioral cloning agent that takes voxel grids as input and outputs 6-DoF actions. While PerAct achieved state-of-the-art performance on RLBench, it has a high computational cost due to processing voxels. Follow-up works, such as RVT [12] and Act3D [13], have reduced the computational cost of PerAct by avoiding voxel representations but often need multiple views of the scene to achieve optimal performance and may be less interpretable compared to a voxel grid that contains a 3D spatial action map. These prior works have not been applied to bimanual manipulation. In this work, we retain the spatial equivariance benefits of voxel representations but reduce the cost of processing voxels by “zooming” into part of the voxel grid. This is similar to the intent of C2F-ARM [47] and RVT-2 [48], but we use the knowledge in VLMs to determine the most relevant regions in the voxel grid.
@@ 3 Problem Statement
Given access to a pre-trained VLM and expert demonstrations, the objective is to produce a bimanual policy $\pi$ for a variety of language-conditioned manipulation tasks. We assume a flat workspace with two fixed-base robot manipulators, each with a parallel-jaw gripper. A policy $\pi$ controls both arms by producing actions $a_{t}=(a_{t}^{s},a_{t}^{a})$ at each time step $t$, where $a_{t}^{s}$ and $a_{t}^{a}$ follow [39] and refer to the stabilizing and acting arm actions, respectively. For simplicity, we suppress the time $t$ when the distinction is unnecessary. We use the low-level action representation from PerAct [11] with $a^{s}=(a_{\rm pose}^{s},a_{\rm open}^{s},a_{\rm collide}^{s})$ and $a^{a}=(a_{\rm pose}^{a},a_{\rm open}^{a},a_{\rm collide}^{a})$. These specify each arm’s 6-DOF gripper pose, its gripper open state, and whether a motion planner for the arms used collision avoidance to reach an intermediate pose. We assume task-specific demonstrations $\mathcal{D}_{\mathcal{\ell}}=\{\zeta_{1},\zeta_{2},\ldots,\zeta_{n}\}$ and two common language commands $\ell_{as}$ and $\ell_{sa}$, where $as$ denotes the left arm as acting and right arm as stabilizing, and vice versa for $sa$. Each demonstration consists of a set of keyframes extracted from a sequence of continuous actions paired with observations. We adapt the keyframe extraction function from [11] by including keyframes that have an action with near-zero joint velocities and unchanged gripper open state for acting and stabilizing arms. The observation at each time is the 3D voxel grid $\mathbf{v}$ of dimension $(L\times W\times H)$, where we use $\mathbf{v}[x,y,z]$ to denote an individual voxel at coordinates $(x,y,z)$. The voxel grid is reconstructed from RGB-D sensors. The robot also receives the language command $\mathbf{l}\in\{\ell_{as},\ell_{sa}\}$, which is fixed for all time steps in an episode, where the robot interacts with the environment for up to $T$ time steps. An episode terminates with a task-dependent success criteria or failure (if otherwise).
@@ 4.1 Extending PerAct for Bimanual Manipulation
PerAct [11] was originally designed and tested for single-arm manipulation. We extend it to support bimanual manipulation. A natural way to do this would be to train separate policies for the two arms. However, we exploit the discretized action space that predicts the next best voxel with spatial equivariance properties and formulate a system that uses acting and stabilizing policies. In contrast to a policy that operates in joint-space control, acting and stabilizing policies perform the same functions irrespective of whether it is a left arm or a right arm, assuming the next best voxel is kinematically feasible for both arms. This policy formulation enables more efficient learning from multi-modal demonstrations compared to a joint-space control policy. In the low-level action space, the arms execute one low-level action $a_{t}^{s}$ and $a_{t}^{a}$ (see Section 3) at each time $t$. In the following, we use similar notation as [11] but index components as belonging to an arm using the superscript: $\textrm{arm}\in\{\textrm{acting},\textrm{stabilizing}\}$.
At each time step, the input to each arm is a voxel observation $\mathbf{v}$, proprioception data of both robot arms $\rho$, a language goal $\mathbf{l}\in\{\ell_{as},\ell_{sa}\}$, and an arm ID $\xi\in\{0,1\}$, and the task is to predict an action. During training, the language goal is given in the data, but during evaluation, we use VLMs to determine which language goal, $\ell_{as}$ or $\ell_{sa}$, to use based on the given task. If the language goal is $\ell_{as}$, we assign the left arm ($\xi=0$) to the acting policy and the right arm ($\xi=1$) to the stabilizing policy, and conversely for $\ell_{sa}$. This allows our method to learn to map the appropriate acting or stabilizing actions to a given arm during training. Note that the predicted arm ID is discarded during evaluation. PerAct uses value maps to represent different components of the action space, where predictions for each arm are $\mathcal{Q}$-functions with state-action values. Formally, we have the following five value maps per arm, as the output of the arm’s learned deep neural network, where:
@@
The demonstrations provide labels for each arm’s five action components, giving us the following nine label sources: ${Y_{\textrm{trans}}^{\textrm{arm}}\in\mathbb{R}^{L\times W\times H}}$ for translations, ${Y_{\textrm{rot}}^{\textrm{arm}}\in\mathbb{R}^{(360/R)\times 3}}$ (with $R=5$ or 5-degree bins) for discretized rotations, ${Y_{\textrm{open}}^{\textrm{arm}}\in\mathbb{R}^{2}}$ for the binary open variables, ${Y_{\textrm{collide}}^{\textrm{arm}}\in\mathbb{R}^{2}}$ for the binary collide variables, and ${Y_{\textrm{id}}^{\textrm{arm}}\in\mathbb{R}^{2}}$ for the binary arm ID variables. The overall training loss for VoxAct-B is:
@@ 4.2 VoxAct-B: Voxel Representations and PerAct for Bimanual Manipulation
To detect the object of interest reliably, we use a two-stage approach similar to [16]. We input a text query and a RGB image from the front camera to OWL-ViT [64], an open-vocabulary object detector, to detect the object. Then, we use Segment Anything [65], a foundational image segmentation model, to obtain the segmentation mask of the object and use the mask’s centroid along with point cloud data, obtained from the front camera’s RGB-D image, to retrieve the object’s pose with respect to the front camera. We use the pose of the object to determine the task-specific roles of each arm and the language goal. This cropped voxel grid and language goal are the input to our bimanual manipulation policy. We call our method VoxAct-B: Voxel-Based Acting and Stabilizing Policy. See Figures 2 and 3 for an overview. Additional implementation details can be found in Appendix A.1.
@@ 5 Experiments
For simulation experiments, we build on top of RLBench [14], a popular robot manipulation benchmark widely used in prior work, including VoxPoser and PerAct. We extend it to support bimanual manipulation and introduce additional variations (see Appendix A.2 for details), making the tasks more challenging. We design the following four bimanual tasks:
  • Open Jar: a jar with a screw-on lid is randomly spawned and scaled from 90% to 100% of the original size within the robot’s $0.43\times 0.48$ meters of workspace. The robot must grasp the jar with one hand and use the other to unscrew the lid in an anti-clockwise direction until it is removed.
@@
  • Hand Over Item: a block is randomly spawned in a workspace of $0.43\times 0.48$ meters. The robot needs to grasp a block with one hand and hand it over to the other.
See Figure 1 for an illustration. In the real world, we test Open Jar and Open Drawer using a coffee jar with dimensions $3.35\times 2.85\times 4.8$ inches and a drawer of dimensions $12\times 12\times 12$ inches. Note that the real-world jar and drawer cannot be opened without the use of a second arm.
@@ 5.1 Baselines and Ablations
  • VoxAct-B w/o Segment Anything: uses the bounding box obtained from OWL-ViT to compute the object’s centroid.
@@
  • VoxAct-B w/o arm ID: disables the arm ID loss function.
@@ 5.2 Experiment Protocol and Evaluation
To generate demonstrations in simulation, we follow the convention from RLBench and define a sequence of waypoints to complete the task, and use motion planning to control the robot arms to reach waypoints. We generate 10 and 100 demonstrations of training data. Half of this data consists of left-acting and right-stabilizing demonstrations, and the other half contains right-acting and left-stabilizing demonstrations. We generate 25 episodes of validation and test data using different random seeds. We train and evaluate all methods using five random seeds and average the results. We evaluate all methods on the same set of test demonstrations for a fair comparison. See Appendix A.1 for details on how checkpoint selection is done in VoxAct-B.
@@ 6.1 Simulation Results
Open Open Put Item Hand Over
Jar Drawer in Drawer Item
Method 10 100 10 100 10 100 10 100
Diffusion Policy 4.8 21.6 4.8 5.6 2.4 4.8 0.0 0.0
ACT w/Transformers 4.0 30.4 12.8 28.0 8.8 44.8 1.6 7.2
VoxPoser 8.0 8.0 32.0 32.0 4.0 4.0 0.0 0.0
Bimanual PerActs 8.0 - 36.8 - 5.6 - 0.0 -
VoxAct-B (ours) 40.0 59.2 73.6 72.8 39.2 49.6 19.2 14.4
Table 1: Performance of different methods on bimanual manipulation tasks in simulation, based on 10 or 100 (task-specific) training demonstrations. We use five training seeds for all methods, and evaluate on the same 25 episodes of unseen test data using the best checkpoints from validation (Section 5.2). The results are the average evaluation over five seeds. We only test Bimanual Peracts with ten demonstrations (not 100) due to computational constraints. VoxPoser does not have training, so its 10 and 100 results are identical.
Comparisons with baselines. Table 1 reports the test success rates of baselines and VoxAct-B. When we train all methods using ten demonstrations, VoxAct-B outperforms all baselines by a large margin. In a low-data regime, the discretized action space with spatial equivariance properties (as used in VoxAct-B and Bimanual PerActs) may be more sample-efficient and easier for learning-based methods compared to methods that use joint space (ACT and Diffusion Policy). When we train all methods using more demonstrations (100), VoxAct-B still outperforms all baselines. Through ablations of ACT and Diffusion Policy, we found that removing environment variations greatly improved their performance. We attribute the tasks’ difficulty to the following: high environment variation, difficult bimanual manipulation tasks with high-dimensional action spaces and fine-grained manipulation, and the two types of training data that each method needs to learn (based on which arms are acting and stabilizing).
@@
VoxAct-B w/o VLMs 19.2
VoxAct-B w/o Segment Anything 67.2
VoxAct-B w/o acting and stabilizing 64.8
@@
Ablation experiments. Table 2 reports results on Open Drawer in simulation, based on 10 training demonstrations and evaluated across five training seeds. We use the same training and evaluation protocols as Table 1. VoxAct-B w/o VLMs performs poorly versus VoxAct-B because, without using the VLMs to reduce the physical space, the voxel resolution is lower due to the large workspace area for each individual voxel, which hinders fine-grained manipulation. VoxAct-B w/o Segment Anything performs worse than VoxAct-B because the predicted object centroid from Segment Anything is closer to the ground truth centroid obtained from the simulator, which is used for training VoxAct-B, than OWL-ViT’s. We found that the closer the predicted object centroid is to the ground truth, the better the policy performs. Moreover, VoxAct-B w/o acting and stabilizing and VoxAct-B w/o arm ID perform worse than VoxAct-B, and they struggle with the same issues as the baselines. See Appendix A for additional experiments and details.